OpenStudy (anonymous):

An equation which represents the average weight in pounds of an adult male of a given height in inches is w = 4.5h - 142. Predict the weight in pounds of an average male with a height of 6' 3" (75 inches). Round to the nearest pound. A. 168 B. 182 C. 196 D. 212

OpenStudy (anonymous):

TAKE HELP FrOM THESE EXAMPLEsss 1, Predit the total cost if you bought 20 books. Books Bought 1 3 5 7 Total Amount Spent $ 21 39 57 75 A, $192 B, $420 C, $200 D, $310 2, An equation which represents the average weight in pounds of an adult male of a given height in inches is w = 4.5h - 142. Predict the weight in pounds of an average male with a height of 6' 3" (75 inches). Round to the nearest pound. A, 168 B, 182 C, 196 D, 212 3, Find the slope for the values in the table below. x -1 0 1 2 y 5 3 1 -1 A, 1 B, -1 C, 2 D, -2 4, Write an equation for the table. x -1 0 1 2 y 5 3 1 -1 A, y = -2x + 3 B, y = -x + 3 C, y = 2x + 5 D, y = x + 5 5, Predict the y-value if x = -3. x -1 0 1 2 y 5 3 1 -1 A, 9 B, -3 C, 3 D, 7 1 year ago Report Abuse 1) A, each old number of books cost $18. From 7-19 is 6 if you count 7.9.11.13.......6 times 18 + 8, because 19-20 is only 1 book, which is 8 dollars because if 2 books cost 18, one cost 9. add them you get 192. 2)C, 4.5+75-142=195.5 round to 196. 3) If 4 and 5 are right, then 3 should be D. The slope is the number before (x) in y=mx+b 4)A, plug in number and it works. 5)A, plug in number, -2(-3)+3=9 1 year ago

OpenStudy (anonymous):

All you do is plug in 75 inches into your weight equation

OpenStudy (anonymous):

tell me what you get, and i'll tell you if it's right

OpenStudy (anonymous):

4.5*(75)-142

Directrix (directrix):

w = 4.5 (75) - 142 w = 4.5* (75) -142 w = 195.5 or 196 to the nearest pound --> C. 196